# From Alignment to Advancement: Bootstrapping Audio-Language Alignment with Synthetic Data

### v2

@@ Abstract
Audio-aware large language models (ALLMs) have recently made great strides in understanding and processing audio inputs. These models are typically adapted from text-based large language models (LLMs) through additional training on audio-related tasks. However, this adaptation process presents two major limitations. First, ALLMs often suffer from catastrophic forgetting, where crucial textual capabilities like instruction-following are lost after training on audio data. In some cases, models may even hallucinate sounds that are not present in the input audio, raising concerns about reliability. Second, achieving cross-modal alignment between audio and language typically relies on large collections of task-specific question–answer pairs for instruction tuning, making it resource-intensive. To address these issues, previous works have leveraged the backbone LLMs to synthesize general-purpose, caption-style alignment data. In this paper, we propose a data generation framework that produces contrastive-like training data, designed to enhance ALLMs’ ability to differentiate between present and absent sounds. We further extend our approach to multi-audio scenarios, enabling the model to either explain differences between audio inputs or produce unified captions that describe all inputs, thereby enhancing audio-language alignment. We refer to the entire ALLM training framework as bootstrapping audio-language alignment via synthetic data generation from backbone LLMs (BALSa). Experimental results indicate that our method effectively mitigates audio hallucinations while reliably maintaining strong performance on audio understanding and reasoning benchmarks, as well as instruction-following skills. Moreover, incorporating multi-audio training further enhances the model’s comprehension and reasoning capabilities. Overall, BALSa offers an efficient and scalable approach to developing ALLMs.
@@ I Introduction
In the audio-language domain, audio-aware large language models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] (ALLMs) refer to models that integrate audio processing capabilities into LLMs, enabling them to understand and generate responses based on auditory inputs. A critical step in developing these models is cross-modal alignment, where textual and auditory information are integrated to ensure effective reasoning across modalities.
To achieve such alignment in practice, the dominant approach relies on instruction tuning with large-scale audio-language datasets [1, 2, 3, 4, 5], typically composed of task-specific question-answer pairs. However, it is expensive to manually collect this kind of data, so some previous work [1, 2, 4, 5, 3] often turns to proprietary language models like ChatGPT [11] to synthesize it using metadata from existing audio corpora.
@@
Beyond the data curation effort, another fundamental issue with existing ALLMs is catastrophic forgetting. Since most ALLMs are adapted from text-based LLMs, they inherently possess strong language understanding capabilities before being trained on audio tasks. However, prior research has reported that currently developed ALLMs often lose their textual capabilities after being trained on audio data. This includes a decline in instruction-following behavior [12], overfitting to pre-defined tasks [3], and limited generalization to unseen cross-modal reasoning tasks [3]. In some cases, the models may even mistakenly identify sounds that are not present in the audio [13].
@@
To address both the challenge of catastrophic forgetting and the burden of data curation, there is a growing trend toward leveraging a backbone LLM to generate synthetic data for audio-language alignment [14, 15, 16, 17, 18, 19]. Here, the backbone LLM refers to the large language model that serves as the initialisation of the core component of ALLM architectures. Instead of relying on manually crafted or task-specific question-answer pairs, which are often generated using proprietary LLMs and complex rules, these approaches leverage the backbone LLM to synthesize training data in a prompt-driven, automatic, and extensible manner. The developers only need to design simple generation prompts, and the backbone LLM can flexibly generate diverse data with minimal human effort. While the conceptual direction of using backbone LLMs for data generation has been widely explored in the speech domain, this article is the first to focus on extending the idea to general audio tasks, demonstrating that it also offers advantages in simplicity, flexibility, and scalability in the audio domain.
@@
This paper represents the framework, BALSa, short for Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs. Different from previous works using backbone LLMs for synthetic data generation, BALSa generates not only descriptive data, such as natural language descriptions of events occurring in the audio, but also contrastive data, which distinguish sounds that are present from those that are absent by simply varying the generation prompt [^1] The incorporation of the synthetic negative samples was first proposed in our previous conference paper, LISTEN (Learning to Identify Sounds Through Extended Negative Samples) [20]. This article extends that conference paper with a unified training framework, multi-audio modeling, and additional evaluations and experiments.. Like previous work of synthetic data generation, BALSa leverages the backbone LLM’s understanding of audio metadata, such as sound event tags, to synthesize training data. Consequently, we only need the prompts like Repeat the audio or List non-existent sounds in the audio to efficiently generate diverse audio-language data tailored to various learning objectives. BALSa offers a flexible and generalizable pipeline that facilitates audio-language alignment by learning from the behavioral patterns of the backbone LLM. This highlights BALSa’s potential as a general-purpose framework for synthesizing diverse alignment data for a wide range of audio understanding tasks. Building upon the flexibility of BALSa in generating diverse alignment data, our findings indicate that incorporating negative synthetic samples effectively mitigates hallucinations, thereby enhancing model reliability.
Furthermore, BALSa can be adapted and extended to various application scenarios. In this study, we introduce an extension of BALSa designed for multi-audio scenarios, where the backbone LLM independently generates explanations of the differences between multiple audio samples and a caption that describes both audio samples together, serving as training data. The objective is to expand the original alignment process into a twofold approach. First, a discrimination-based process that generates comparative explanations for differences between audio samples. Second, a description-based process that generates separate captions for each audio sample within the same response.
Our experimental results demonstrate that training the model on multi-audio data, whether through joint captioning or comparative discrimination, sufficiently enhances its overall performance. Interestingly, the difference between generating a single caption that encompasses both audio samples and producing comparative explanations for their differences is minimal. This suggests that simply learning to handle multiple audio samples at once is sufficient to improve the model’s comprehension, even without explicit comparative training. These findings indicate that exposing the model to multiple audio inputs, whether for captioning (descriptive learning) or discrimination (comparative learning), is a key factor in boosting performance. Furthermore, the BALSa-based approach leads to more efficient data usage. Our method achieves comparable or even better performance while using only 12% of the training data duration required by the previous state-of-the-art model [2].
@@
    Prior works [14, 15, 16, 17, 18, 19] have explored and validated the effectiveness of synthetic data generation using backbone LLMs. In these studies, the backbone LLM analyzes audio metadata and generates descriptions guided by designed prompts, but they focus primarily on the speech domain. This article is the first to apply this paradigm specifically to the non-speech audio domain. BALSa synthesizes audio-language alignment data using simple prompts from audio metadata, such as sound event tags, without requiring task-specific question-answer collections.
@@
    BALSa demonstrates high data efficiency, achieving performance comparable to or better than a widely adopted strong model [2], while using only 12% of the training data duration.
@@
    To comprehensively assess audio-aware large language models, we conduct experiments on existing benchmarks, reproduce missing evaluation sets, and introduce new test sets to explore previously unexamined aspects. Our evaluation covers three key areas: audio question answering, which assesses fundamental audio understanding; audio reasoning; and reliability and safety. Through these evaluations, we systematically compare BALSa with mainstream models, examining the effectiveness of audio-language alignment and addressing challenges like audio hallucinations and instruction adherence.
@@ II Related Works
There is already extensive research on audio and speech-related large language models. Due to space constraints, this paper discusses only the studies directly relevant to our work. For a comprehensive overview of the broader framework of audio and speech LLMs, readers are encouraged to consult the survey papers [21, 22, 23, 24, 25, 26, 27, 28].
Large language models (LLMs) [29, 30, 31, 32, 33] can perform zero-shot learning through instruction tuning. This capability is expected to emerge when strong cross-modal alignment is achieved, particularly in scenarios where a text-based LLM is integrated with multi-modal encoders. Following this trend, audio-aware large language models [1, 2, 3, 4, 5, 6, 7, 8, 34] extend instruction tuning to audio-related tasks by combining audio encoders with language models and training them in a multi-stage process.
@@
On the other hand, a distinct strategy for developing speech- or audio-aware large language models focuses on leveraging the backbone LLM itself to generate synthetic data for alignment. These alignment data include general response-style texts conditioned on speech/audio, captions describing speech/audio under textual conditions, or continuation responses of speech/audio content based on textual prompts. Several previous works [16, 14, 17, 19, 18] have adopted this approach. Unlike studies that rely on externally generated task-specific question-answer pairs [1, 2, 3, 4, 5], these works utilized a backbone language model to generate general-purpose alignment data directly, enabling a more streamlined and self-contained approach to integrating speech/audio with language models. DeSTA [15] found that the approach leveraged speech captioning to bridge the gap between speech and text modalities, while DeSTA2 [14] further discovered that generating captions using the backbone LLM was more efficient. AudioChatLlama [16] and DiVA [19] used the backbone language model to generate responses conditioned on textual transcriptions. BLSP [17] further extended this idea by generating continuations based on transcriptions, while BLSP-Emo [18] incorporated additional paralinguistic information like emotion to produce more expressive outputs. DeSTA2 [14] further extended the use of paralinguistic cues, such as emotion, accent, and gender, to enhance the expressiveness of generated speech captions.
Inspired by these methods and motivated by the lack of research applying this strategy to general audio settings, our work investigates how to bootstrap audio–language alignment through synthetic data generation using a backbone LLM. We further explore various extensions of this idea to evaluate its generality and effectiveness across different scenarios. For example, we exploit the flexibility and scalability of this framework to introduce negative samples and multi-audio scenario into the training process. This article provides a practical way to effectively train audio-aware LLMs without the need to manually collect or rely on proprietary models to generate a large amount of task-specific question-answer data.
@@ III-A Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs
The rationale behind generating audio-language pairs using the backbone LLM is to enable the model to acquire audio understanding capabilities solely through an audio modality adapter. This approach ensures that the audio-aware large language model aligns its comprehension of audio with the text-based metadata information understood by the backbone LLM, without requiring adjustments to the backbone LLM itself.
@@
Positive Samples Generation Prompt (${P}_{{pos}}$) aims to generate descriptions of sound events that actually occur in the audio. For example, Replay the audio.
@@ III-B BALSa-MA: Bootstrapping Multi-Audio Learning via Difference Comparison and Joint Captioning
Explain the difference between two audio.
@@
Repeat the both audio.
@@
Audio1: The sound of water pouring accompanies a woman speaking nearby.
Audio2: A man talks while a dog barks in the background.
@@ III-C Model Design and Training Approach
Backbone Large Language Model. This study adopts LLaMA-3.1-8B-Instruct [30], an instruction-tuned model, as the backbone large language model. To maintain its original text-processing capabilities, we do not apply LoRA or fine-tune any of its parameters.
Audio Modality Adapter. The only trainable component is the audio modality adapter, which is randomly initialized. This adapter projects the output representations extracted by the audio encoder into the input space of the backbone large language model. Specifically, we employ a Qformer [39] architecture to extract audio features from the intermediate layers of the Whisper encoder. These extracted features are aggregated using a weighted sum with learnable weights, followed by a linear projection layer that aligns the aggregated features with the input dimensions of the backbone large language model. We optimize the entire architecture end-to-end with a next-token prediction loss, using the model’s own generated responses as learning targets. To improve efficiency, we freeze the parameters of both the backbone LLM and the audio encoder. Instead, we adopt a lightweight adapter-based training strategy [40, 15, 17, 18, 19], in which only the modality adapter is trained. This design not only conserves computational resources but also preserves the LLM’s original linguistic knowledge while enabling it to process audio through the adapter.
@@ IV-A Training Datasets
The training datasets include AudioSet-20K [41], AudioCaps [42], FSD50K [43], MACS [44], ESC-50 [45], UrbanSound8K [46], Clotho [47], and VocalSound [48]. Among these datasets, AudioCaps, Clotho, and MACS provide ground-truth labels in the form of human-annotated audio captions, while the rest use sound event tags. Table III summarizes the total duration of the training datasets used in our work and compares it with the training data durations reported in previous studies [1, 2, 4, 5, 3]. The “Duration” column represents the original total duration of unique audio in each dataset, measured in hours and excluding any duplicate segments. The “Equi. Duration” column accounts for the effect of generating multiple aligned text descriptions per audio clip, leading to an increased effective training duration, also measured in hours. The blue numbers indicate the amount of data used in the multi-audio scenario. The number of samples represents the total number of data instances used during training.
@@
Dataset Duration Equi. Duration # Samples
AudioSet-20K[41] 60 121 42K
AudioCaps[42] 138 276+269 96K+49K
Clotho[47] 18 36+160 27K+13K
MACS[44] 11 22 33K
FSD50K[43] 56 113+75 56K+19K
ESC50[45] 3 6+28 4K+10K
UrbanSound8K[46] 10 19 17K
VocalSound[48] 20 40 34K
Total (Ours) 316 1164 400K
Qwen2-Audio[1, 2] - 10000 -
SALMONN[3] - 1044 370K
LTU[4], LTU-AS[5] - 1784 5682K

### v3

@@ Abstract
Audio-aware large language models (ALLMs) have recently made great strides in understanding and processing audio inputs. These models are typically adapted from text-based large language models (LLMs) through additional training on audio-related tasks. This adaptation process presents two major limitations. First, ALLMs often suffer from catastrophic forgetting, where crucial textual capabilities like instruction-following are lost after training on audio data. In some cases, models may even hallucinate sounds that are not present in the input audio, raising concerns about reliability. Second, achieving cross-modal alignment between audio and language typically relies on large collections of task-specific question–answer pairs for instruction tuning, making it resource-intensive. To address these issues, previous works have leveraged the backbone LLMs to synthesize general-purpose, caption-style alignment data. In this paper, we propose a data generation framework that produces contrastive-like training data, designed to enhance ALLMs’ ability to differentiate between present and absent sounds. We further extend our approach to multi-audio scenarios, enabling the model to either explain differences between audio inputs or produce unified captions that describe all inputs, thereby enhancing audio-language alignment. We refer to the entire ALLM training framework as bootstrapping audio-language alignment via synthetic data generation from backbone LLMs (BALSa). Experimental results indicate that our method effectively mitigates audio hallucinations while reliably maintaining strong performance on audio understanding and reasoning benchmarks, as well as instruction-following skills. Moreover, incorporating multi-audio training further enhances the model’s comprehension and reasoning capabilities. Overall, BALSa offers an efficient and scalable approach to developing ALLMs.
@@ I Introduction
In the audio-language domain, audio-aware large language models [5, 6, 56, 26, 25, 21, 36, 11, 38] (ALLMs) refer to models that integrate audio processing capabilities into LLMs, enabling them to understand and generate responses based on auditory inputs. A critical step in developing these models is cross-modal alignment, where textual and auditory information are integrated to ensure effective reasoning across modalities.
@@
To achieve such alignment in practice, the dominant approach relies on instruction tuning with large-scale audio-language datasets [5, 6, 56, 26, 25], typically composed of task-specific question-answer pairs. However, it is expensive to manually collect this kind of data, so some previous work [5, 6, 26, 25, 56] often turns to proprietary language models like ChatGPT [2] to synthesize it using metadata from existing audio corpora. Beyond the data curation effort, another fundamental issue with existing ALLMs is catastrophic forgetting. Since most ALLMs are adapted from text-based LLMs, they inherently possess strong language understanding capabilities before being trained on audio tasks. However, prior research has reported that currently developed ALLMs often lose their textual capabilities after being trained on audio data. This includes a decline in instruction-following behavior [43], overfitting to pre-defined tasks [56], and limited generalization to unseen cross-modal reasoning tasks [56]. In some cases, the models may even mistakenly identify sounds that are not present in the audio [37].
@@
To address both the challenge of catastrophic forgetting and the burden of data curation, there is a growing trend toward leveraging a backbone LLM to generate synthetic data for audio-language alignment [16, 59, 60, 29]. Here, the backbone LLM refers to the large language model that serves as the initialisation of the core component of ALLM architectures. Instead of relying on manually crafted or task-specific question-answer pairs, which are often generated using proprietary LLMs and complex rules, these approaches leverage the backbone LLM to synthesize training data in a prompt-driven, automatic, and extensible manner. The developers only need to design simple generation prompts, and the backbone LLM can flexibly generate diverse data with minimal human effort. While foundational datasets like Audio Dialogues [22], CoTA [64], and CompA [19] have been instrumental, their creation methodologies differ fundamentally from an emerging paradigm: leveraging the backbone LLM itself to synthesize training data [16, 59, 60, 29]. While this approach has been widely explored in the speech domain, it has not yet been extended to general audio. This paper is the first to bridge this gap, demonstrating the advantages of this paradigm in simplicity, flexibility, and scalability for diverse general audio tasks.
@@
This paper represents the framework, BALSa, short for Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs. Different from previous works [44, 16, 59, 60, 29] using backbone LLMs for synthetic data generation, BALSa generates not only descriptive data, such as natural language descriptions of events occurring in the audio, but also contrastive data, which distinguish sounds that are present from those that are absent by simply varying the generation prompt [^1] The incorporation of the synthetic negative samples was first proposed in our previous conference paper, LISTEN (Learning to Identify Sounds Through Extended Negative Samples) [39]. This article extends that conference paper with a unified training framework, multi-audio modeling, and additional evaluations and experiments.. Like previous work of synthetic data generation, BALSa leverages the backbone LLM’s understanding of audio metadata, such as sound event tags, to synthesize training data. Consequently, we only need the prompts like Repeat the audio or List non-existent sounds in the audio to efficiently generate diverse audio-language data tailored to various learning objectives. BALSa offers a flexible and generalizable pipeline that facilitates audio-language alignment by learning from the behavioral patterns of the backbone LLM. This highlights BALSa’s potential as a general-purpose framework for synthesizing diverse alignment data for a wide range of audio understanding tasks.
In summary, this work introduces BALSa, a framework with several novel contributions that distinguish it from prior works. First, we are the first to adapt the paradigm of backbone-LLM-driven synthetic data generation from the speech domain to general audio tasks, which cover a diverse range of non-speech events. This methodological shift allows for the generation of varied data formats tailored to specific learning objectives, such as description, comparison, or discrimination, simply by altering the generation prompts. Second, we leverage this capability to address a critical challenge in the field: audio hallucination. Our proposed contrastive-style data generation, particularly the inclusion of synthetic negative samples, directly trains the model to differentiate present from absent sounds, effectively enhancing its reliability. Finally, we demonstrate the scalability and efficiency of our framework. We extend it to multi-audio scenarios (BALSa-MA) to enable more complex reasoning, such as comparative analysis and joint captioning. Furthermore, our approach is highly data-efficient, achieving competitive results while using only a fraction of the training data required by previous state-of-the-art models.
@@
    Prior works [44, 16, 59, 60, 29] have explored and validated the effectiveness of synthetic data generation using backbone LLMs. In these studies, the backbone LLM analyzes audio metadata and generates descriptions guided by designed prompts, but they focus primarily on the speech domain. This article is the first to apply this paradigm specifically to the general audio domain. BALSa synthesizes audio-language alignment data using simple prompts from audio metadata, such as sound event tags, without requiring task-specific question-answer collections.
@@
    The BALSa framework achieves competitive performance with recent state-of-the-art models across reasoning, understanding, and trustworthiness benchmarks, while being data-efficient and even surpassing some models trained with substantially larger datasets.
@@
    To comprehensively assess ALLMs, we conduct experiments on existing benchmarks, reproduce missing evaluation sets, and introduce new test sets to explore previously unexamined aspects. Our evaluation covers three key areas: audio question answering, which assesses fundamental audio understanding; audio reasoning; and reliability and safety. Through these evaluations, we systematically compare BALSa with mainstream models, examining the effectiveness of audio-language alignment and addressing challenges like audio hallucinations and instruction adherence.
  6. 6.
    We further perform complementary analyses to examine the robustness and generalizability of BALSa. These include backbone replacement (Qwen3-4B vs. LLaMA-3.1-8B), LoRA-based fine-tuning, prompt design variations in synthetic data generation, and data scaling experiments. The results show that BALSa remains stable under different backbones and training strategies, achieves more reliable performance when combining diverse prompts, and consistently benefits from larger training data sizes. These findings highlight the framework’s robustness and potential for further improvements at scale.
@@ II Related Works
There is already extensive research on audio and speech-related large language models. Due to space constraints, this paper discusses only the studies directly relevant to our work. For a comprehensive overview of the broader framework of audio and speech large language models (LLMs), readers are encouraged to consult the survey papers [3, 34, 50, 10, 63, 28, 48, 69].
LLMs [58, 15, 49, 66, 70] can perform zero-shot learning through instruction tuning, which emerges when strong cross-modal alignment is achieved, particularly when a text-based LLM is integrated with multimodal encoders. Following this trend, audio-aware large language models (ALLMs) [5, 6, 56, 26, 25, 21, 36, 11, 13, 9, 65, 1, 64, 23, 20, 22, 45, 14, 30, 62] extend this paradigm to the audio domain by combining audio encoders with language models in a multi-stage training process.
@@
Recent models [21, 20, 23, 65, 64] further advance this multi-stage paradigm by strengthening the audio encoder, scaling data, or enhancing reasoning. Audio Flamingo 3 [23] introduces a unified AF-Whisper encoder and large curated datasets (e.g., AudioSkills-XL, AF-Think) with a curriculum strategy, enabling long-audio understanding, multi-audio dialogue, and on-demand chain-of-thought reasoning. Qwen2.5-Omni [65] extends beyond audio to an end-to-end multimodal system, adopting a Thinker–Talker architecture for comprehension and real-time speech generation. Audio-Reasoner [64] builds on Qwen2-Audio but focuses on explicit reasoning, leveraging a 1.2M chain-of-thought dataset (CoTA) to yield large performance gains on reasoning benchmarks.
On the other hand, a distinct strategy for developing speech- or audio-aware LLMs focuses on leveraging the backbone LLM itself to generate synthetic data for alignment. These alignment data include general response-style texts conditioned on speech/audio, captions describing speech/audio under textual conditions, or continuation responses of speech/audio content based on textual prompts. Several previous works [16, 44, 59, 29, 60] have adopted this approach. Unlike studies that rely on externally generated task-specific question-answer pairs [5, 6, 56, 26, 25], these works utilized a backbone language model to generate general-purpose alignment data directly, enabling a more streamlined and self-contained approach to integrating speech/audio with language models. AudioChatLlama [16] and DiVA [29] used the backbone language model to generate responses conditioned on textual transcriptions. BLSP [59] further extended this idea by generating continuations based on transcriptions, while BLSP-Emo [60] incorporated additional paralinguistic information like emotion to produce more expressive outputs. DeSTA2 [44] further extended the use of paralinguistic cues to enhance the expressiveness of generated speech captions.
Inspired by these methods and motivated by the lack of research applying this strategy to general audio settings, our work investigates how to bootstrap audio–language alignment through synthetic data generation using a backbone LLM. We further explore various extensions of this idea to evaluate its generality and effectiveness across different scenarios. For example, we exploit the flexibility and scalability of this framework to introduce negative samples and multi-audio scenario into the training process. This article provides a practical way to effectively train ALLMs without the need to manually collect or rely on proprietary models to generate a large amount of task-specific question-answer data.
@@ III-A Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs
The rationale behind generating audio-language pairs using the backbone LLM is to enable the model to acquire audio understanding capabilities solely through an audio modality adapter. This approach ensures that the ALLM aligns its comprehension of audio with the text-based metadata information understood by the backbone LLM, without requiring adjustments to the backbone LLM itself.
@@
Positive Samples Generation Prompt (${P}_{{pos}}$) aims to generate descriptions of sound events that actually occur in the audio. For example, Can you describe what is in the audio.
@@
In summary, the final input prompt (${P}_{{final}}$) to the LLM can be expressed in Equation 1 and the overall pipeline is illustrated in Figure 1. We intentionally adopt simple and direct prompts to establish a strong baseline, avoiding reliance on complex prompt engineering.
@@ III-B BALSa-MA: Bootstrapping Multi-Audio Learning via Difference Comparison and Joint Captioning
$P_{final,1}$ $=[\textit{Begin of audio1}]\,P_{seed,1}\,[\textit{End of audio1}]$ (2)
$P_{final,2}$ $=[\textit{Begin of audio2}]\,P_{seed,2}\,[\textit{End of audio2}]$
@@
Explain the difference between the two audios.
@@
Describe both audios.
@@
Audio1: The sound of water pouring accompanies a woman speaking nearby. Audio2: A man talks while a dog barks in the background.
@@ III-C Model Design and Training Approach
Backbone Large Language Model. This study adopts LLaMA-3.1-8B-Instruct [15] and Qwen3-4B, as the backbone large language models. To maintain its original text-processing capabilities, we do not fine-tune any of its parameters.
Audio Modality Adapter. The only trainable component is the audio modality adapter, which is randomly initialized. This adapter projects the output representations extracted by the audio encoder into the input space of the backbone large language model. Specifically, we employ a Qformer [40] architecture to extract audio features from the intermediate layers of the Whisper encoder. These extracted features are aggregated using a weighted sum with learnable weights, followed by a linear projection layer that aligns the aggregated features with the input dimensions of the backbone large language model. We optimize the entire architecture end-to-end with a next-token prediction loss, using the model’s own generated responses as learning targets. To improve efficiency, we freeze the parameters of both the backbone LLM and the audio encoder. Instead, we adopt a lightweight adapter-based training strategy [71, 59, 60, 29], in which only the modality adapter is trained. This design not only conserves computational resources but also preserves the LLM’s original linguistic knowledge while enabling it to process audio through the adapter.
@@ IV-A Training Datasets
The training datasets include AudioSet-20K [18], AudioCaps [35], FSD50K [17], MACS [47], ESC-50 [51], UrbanSound8K [55], Clotho [42], and VocalSound [24]. Among these datasets, AudioCaps, Clotho, and MACS provide ground-truth labels in the form of human-annotated audio captions, while the rest use sound event tags. Table III summarizes the total duration of the training datasets used in our work and compares it with the training data durations reported in previous studies [5, 6, 26, 25, 56]. The “Duration” column represents the original total duration of unique audio in each dataset, measured in hours and excluding any duplicate segments. The “Equi. Duration” column accounts for the effect of generating multiple aligned text descriptions per audio clip, leading to an increased effective training duration, also measured in hours. The number of samples represents the total number of data instances used during training.
@@
Dataset Duration Equi. Duration # Samples (k)
AudioSet-20K [18] 60 121 42
AudioCaps [35] 138 545 145
Clotho [42] 18 196 40
MACS [47] 11 22 33
FSD50K [17] 56 188 75
ESC50 [51] 3 34 14
UrbanSound8K [55] 10 19 17
VocalSound [24] 20 40 34
Total (Ours) 316 1,164 400
Qwen2-Audio [5, 6] - 10,000 -
SALMONN [56] - 1,044 370
LTU [26], LTU-AS[25] - 1,784 5,682
GAMA [21] - - 6,319
Qwen-2.5-Omni [65] - 3,333,333 -
Audio Flamingo 2 [20] - 33,153 1,460
Audio Flamingo 3 [23] - 54,409 2,600
Audio-Reasoner [64] - - 1,200
Phi-4-Multimodal-IT [1] - - 17,000